OpenStudy (loser66):

Express 1.191191119..... in summation notation Please, help

OpenStudy (loser66):

@freckles

OpenStudy (arthur326):

To get started, can you express this as a sum with good-ol'-fashioned "+" signs?

OpenStudy (loser66):

\(a_0= 10^0\\a_1=(10^1+9)*10^{-2}\\ a_2=(10^2+a_1)*10^{-3}\\\cdots\\ a_n=(10^n+a_{n-1})*10^{-(n+1)}\)

OpenStudy (loser66):

But I stuck there.

OpenStudy (arthur326):

Sorry I was having connection issues.

OpenStudy (loser66):

So, if I add the LHS, I got 1.191191119..., but I don't know how to do with the rRHS

OpenStudy (arthur326):

I don't know if recursion is particularly helpful here. Now I'm sure there are many ways to go about this, but here's one I see: \[1.191191191... = 119 \cdot 10^{-2} + 119 + \cdot 10^{-5} + 119 \cdot 10^{-8} +\cdots\]

OpenStudy (arthur326):

That grouping is fairly easy to express with summation notation. Do you see how?

OpenStudy (loser66):

the third one is not that

OpenStudy (arthur326):

Please clarify.

OpenStudy (loser66):

\(119*10^{-8}= 0.00000119\)

OpenStudy (loser66):